OpenStudy (anonymous):

he length of a rectangle is 3 yrd more than twice its width, and the area of the rectangle is 54yrds^3 . Find the dimensions of the rectangle.

OpenStudy (btaylor):

L = 3+2W L x W = 54 (W)(3+2W)=54 3W+2W^2=54 2W^2+3W-54=0 by quadratic formula, W = 4.5, so L = 12